OpenStudy (anonymous):

Calculate the derivative of y=4log(base3)(x^2-1)

OpenStudy (anonymous):

\[differentiate y=\log_{3} (x ^{2}-1)\]

OpenStudy (anonymous):

using log_b(z) = log(z)/log(b) y={4log(x^2-1)}/log3 y'=(4/log3)(2x)/(x^2-1) as derivative of log(x^2-1)=2x/x^2-1 y'=8x/(x^2-1)log3

OpenStudy (anonymous):

im kind of confused can you further help me?

OpenStudy (anonymous):

the change of base formula. log(base)(argument) = log(argument) / log(base) standard-base log of the argument) divided by (same-standard-base log of the non-standard-base) i use this formula than take derivative, as 4/log3 is constant and derivative of log(x^2-1) is 2x/(x^2-1) i hope u understand